OpenStudy (anonymous):

Multiply: 5 the square root of 8 • and 4 the square root of 7

OpenStudy (anonymous):

a. 16 the square root of 10 b. 20 the square root of 15 c. 40 the square root of 14 d. 40

OpenStudy (anonymous):

\[5\sqrt{8}\times4\sqrt{7}\]\[10\sqrt{2}\times4\sqrt{7}\]\[40\sqrt{14}\]